OpenStudy (anonymous):

Find m if the line y=mx−2 intersects y=x2 in just one point.

OpenStudy (anonymous):

Really need help

OpenStudy (anonymous):

Will fan and medal

OpenStudy (anonymous):

Can you help me?

OpenStudy (photon336):

I'm wondering if you can just set those two equal to each-other

OpenStudy (anonymous):

i dont know

OpenStudy (mathmale):

That would help! However, I think you could get much more insight into this problem if you would graph y=x^2 and then, on the same set of axes, for values of m = to 1, 2, 3, -1, and so on, you also graph y=mx-2. Why do this? because you're told that the line intersects the curve ONLY ONCE, and it's easiest to determine an appropriate m value visuallyl.

OpenStudy (anonymous):

can you give me the answer

OpenStudy (mathmale):

No. Sorry. That's completely against OpenStudy's Code of Conduct (which I hope you've read). This is a very interesting problem. I've graphed both y=x^2 and y=mx-2 on my calculator for several different values of m, and am close to identifying a solution in that manner.

OpenStudy (mathmale):

Photon suggested earlier: "I'm wondering if you can just set those two equal to each-other ." Why not try that? Set x^2 = y = y = mx-2 and solve for m. Try it.

OpenStudy (mathmale):

Unfortunately, x^2=mx-2 has TWO variables in it, which means we don't have enuf info to solve for m. So, back to the drawing board. You can estimate m by choosing various m values and graphing both x^2 and mx-2. For approx. which m value will the line just touch the curve in ONE place only?

OpenStudy (mathmale):

I'm sorry, but I have to get off OpenStudy. I still feel your best bet would be the graphical approach I've suggested, although I'm sure there are analytic (algebraic) approaches that will succeed.